OpenStudy (anonymous):

Complete the identity: a) sin(A-B) = b) tan(A+B) = c) cos(2A) = (3 answers) helllllllp

hartnn (hartnn):

do u just need the formula? or u need to derive it ?

OpenStudy (anonymous):

he just said complete...not sure what that means honestly

hartnn (hartnn):

ok, then u just need the formulas: here, sin (A-B) = sin A cos b -cos A sin B

hartnn (hartnn):

tan (A+B) = (tan A+ tan B)/ (1+ tan A .tan B)

hartnn (hartnn):

cos 2A is important. cos 2A= cos^2A-sin ^2 A = 1-2sin^2 A =2cos^2A -1

hartnn (hartnn):

these are idinteties, remember them

OpenStudy (anonymous):

wow thats confusing!! so much stuff. thanks!

hartnn (hartnn):

welcome :)